OpenStudy (anonymous):

Maria and some of her friends are on a road trip. Maria was assigned to drive 3/7 of the total trip. She has already driven3/5 of her assigned driving duty. How much of the total trip has Maria driven?

OpenStudy (tkhunny):

Have you considered multiplication?

OpenStudy (misty1212):

lol HI!!

OpenStudy (misty1212):

three fifth OF 2 sevenths means \[\frac{3}{5}\times \frac{3}{7}\]